you own a portfolio that is invested 25 percent in stock A, 40 percent in stock B, and the reminder in stock C. the expected returns on stocks are 8 percent, 16 percent, and 5 percent respectively. what is the expected return on the portfolio?

a. 8.28%

b. 10.15%

c. 12.39%

d. 13.14%

e. 14.58%
Answer #1

expected return on the portfolio=Respective return*Respective weight

=(0.25*8)+(0.4*16)+(0.35*5)(Investment in C=(100-25-40)=35%)

which is equal to

=10.15%